# Understanding World or Predicting Future? A Comprehensive Survey of World Models

### v1

@@ Abstract
The concept of world models has garnered significant attention due to advancements in multimodal large language models such as GPT-4 and video generation models such as Sora, which are central to the pursuit of artificial general intelligence. This survey offers a comprehensive review of the literature on world models. Generally, world models are regarded as tools for either understanding the present state of the world or predicting its future dynamics. This review presents a systematic categorization of world models, emphasizing two primary functions: (1) constructing internal representations to understand the mechanisms of the world, and (2) predicting future states to simulate and guide decision-making. Initially, we examine the current progress in these two categories. We then explore the application of world models in key domains, including autonomous driving, robotics, and social simulacra, with a focus on how each domain utilizes these aspects. Finally, we outline key challenges and provide insights into potential future research directions.
1 Introduction
@@
Figure 1: The overall framework of this survey. We systematically define the essential purpose of a world model as understanding the dynamics of the external world and predicting future scenarios. The timeline illustrates the development of key definitions and applications.
@@
The scientific community has long aspired to develop a unified model that can replicate its fundamental dynamics of the world in pursuit of Artificial General Intelligence (AGI) [98]. In 2024, the emergence of multimodal large language models (LLMs) and Sora [130] have intensified discussions surrounding such World Models. While these models demonstrate an emerging capacity to capture aspects of world knowledge — such as Sora’s generated videos, which appear to perfectly adhere to physical laws — questions persist regarding whether they truly qualify as comprehensive world models. Therefore, a systematic review of recent advancements, applications, and future directions in world model research is both timely and essential as we look toward new breakthroughs in the era of artificial intelligence.
@@
The definition of a world model remains a subject of ongoing debate, generally divided into two primary perspectives: understanding the world and predicting the future. As depicted in Figure 1, early work by Ha and Schmidhuber [59] focused on abstracting the external world to gain a deep understanding of its underlying mechanisms. In contrast, LeCun [98] argued that a world model should not only perceive and model the real world but also possess the capacity to envision possible future states to inform decision-making. Video generation models such as Sora represent an approach that concentrates on simulating future world evolution and thus align more closely with the predictive aspect of world models. This raises the question of whether a world model should prioritize understanding the present or forecasting future states. In this paper, we provide a comprehensive review of the literature from both perspectives, highlighting key approaches and challenges.
@@
The potential applications of world models span a wide array of fields, each with distinct requirements for understanding and predictive capabilities. In autonomous driving, for example, world models need to perceive road conditions in real-time [195, 177] and accurately predict their evolution [127, 167, 241], with a particular focus on immediate environmental awareness and forecasting of complex trends. For robotics, world models are essential for tasks such as navigation [160], object detection [183], and task planning [62], requiring a precise understanding of external dynamics [47] and the ability to generate interactive and embodied environments [132]. In the realm of simulation of virtual social systems, world models must capture and predict more abstract behavioral dynamics, such as social interactions and human decision-making processes. Thus, a comprehensive review of advancements in these capabilities, alongside an exploration of future research directions and trends, is both timely and essential.

### v2

@@ Abstract.
The concept of world models has garnered significant attention due to advancements in multimodal large language models such as GPT-4 and video generation models such as Sora, which are central to the pursuit of artificial general intelligence. This survey offers a comprehensive review of the literature on world models. Generally, world models are regarded as tools for either understanding the present state of the world or predicting its future dynamics. This review presents a systematic categorization of world models, emphasizing two primary functions: (1) constructing internal representations to understand the mechanisms of the world, and (2) predicting future states to simulate and guide decision-making. Initially, we examine the current progress in these two categories. We then explore the application of world models in key domains, including autonomous driving, robotics, and social simulacra, with a focus on how each domain utilizes these aspects. Finally, we outline key challenges and provide insights into potential future research directions. We summarize the representative papers along with their code repositories in https://github.com/tsinghua-fib-lab/World-Model.
World model, model-based RL, video generation, embodied environment, autonomous driving, robots, social simulacra
*These two authors contributed equally.
†These authors contributed equally.
copyright: acmcopyrightjournalyear: 20xxjournal: CSURcopyright: noneccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Modeling and simulation
1. Introduction
@@
Figure 1. The overall framework of this survey. We systematically define the essential purpose of a world model as understanding the dynamics of the external world and predicting future scenarios. The timeline illustrates the development of key definitions and applications.
@@
The scientific community has long aspired to develop a unified model that can replicate its fundamental dynamics of the world in pursuit of Artificial General Intelligence (AGI) (lecun2022path, ). In 2024, the emergence of multimodal large language models (LLMs) and video generation models like Sora (sora2024, ) has intensified discussions surrounding such World Models. While these models demonstrate an emerging capacity to capture aspects of world knowledge–such as Sora’s generated videos, which appear to perfectly adhere to physical laws–questions persist regarding whether they truly qualify as comprehensive world models. Therefore, a systematic review of recent advancements, applications, and future directions in world model research is both timely and essential as we look toward new breakthroughs in the era of artificial intelligence (AI).
@@
The definition of a world model remains a subject of ongoing debate, generally divided into two primary perspectives: understanding the world and predicting the future. As depicted in Figure 1, early work by Ha and Schmidhuber (ha2018world, ) focused on abstracting the external world to gain a deep understanding of its underlying mechanisms. In contrast, LeCun (lecun2022path, ) argued that a world model should not only perceive and model the real world but also possess the capacity to envision possible future states to inform decision-making. Video generation models such as Sora represent an approach that concentrates on simulating future world evolution and thus align more closely with the predictive aspect of world models. This raises the question of whether a world model should prioritize understanding the present or forecasting future states. In this paper, we provide a comprehensive review of the literature from both perspectives, highlighting key approaches and challenges.
@@
The potential applications of world models span a wide array of fields, each with distinct requirements for understanding and predictive capabilities. In autonomous driving, for example, world models need to perceive road conditions in real-time (yolop_2022, ; teichmann2018multinetrealtimejointsemantic, ) and accurately predict their evolution (ngiam2021scene, ; shi2022motion, ; zhou2023query, ), with a particular focus on immediate environmental awareness and forecasting of complex trends. For robotics, world models are essential for tasks such as navigation (shah2023gnm, ), object detection (wang2024repvit, ), and task planning (hafner2019learning, ), requiring a precise understanding of external dynamics (gao2023s, ) and the ability to generate interactive and embodied environments (park2023generative, ). In the realm of simulation of virtual social systems, world models must capture and predict more abstract behavioral dynamics, such as social interactions and human decision-making processes. Thus, a comprehensive review of advancements in these capabilities, alongside an exploration of future research directions and trends, is both timely and essential.